OpenStudy (anonymous):

please help...Find the measure of exterior angle A. Show equations and all work that leads to your answer.

OpenStudy (anonymous):

how do i solve for x

OpenStudy (phi):

(14x+6)+(6x-3)+(4x+8)+(49)+(3x+30) = 360 How many x's do you have? (Add them up) add up the simple numbers. Can you do that?

OpenStudy (anonymous):

4 x's, & 123

OpenStudy (phi):

14x+6x+4x+3x = how many x's?

OpenStudy (anonymous):

4

OpenStudy (phi):

The number in front of each x is the number of x's 14x is short hand for 14 x's Now, how many x's?

OpenStudy (anonymous):

oh, 27

OpenStudy (phi):

yes, but write it as 27x now add up the numbers 6-3+8+49+30

OpenStudy (anonymous):

90

OpenStudy (phi):

OK, what you just did is simplify (14x+6)+(6x-3)+(4x+8)+(49)+(3x+30) = 360 to 27x+90 = 360 (You added up the terms on the left hand side) Now subtract 90 from both sides of the equation.

OpenStudy (phi):

27x + 90 -90 = 360 -90

OpenStudy (anonymous):

270

OpenStudy (phi):

that is only the right hand side. Write down both sides of the equation

OpenStudy (phi):

if you subtract 90 from both sides you get 27x= 270

OpenStudy (phi):

Now divide BOTH SIDES by 27

OpenStudy (anonymous):

0.1?

OpenStudy (phi):

\[ \frac{27x}{27} = \frac{270}{27} \] can you simplify this?

OpenStudy (anonymous):

no

OpenStudy (phi):

what is 27/27?

OpenStudy (phi):

If you don't know, you can use a calculator

OpenStudy (anonymous):

1

OpenStudy (phi):

That means \[ \frac{27x}{27} = \frac{270}{27} \] is \[ 1x= \frac{270}{27} \] or because we don't usually write a 1, just \[ x= \frac{270}{27} \] now what is 270/27 ?

OpenStudy (anonymous):

10

OpenStudy (phi):

so now you know x=20 Angle A = 6x-3 to find A, replace x with 10 A= 6*10-3 (we have 6 10's then subtract 3) What is A?

OpenStudy (phi):

*x=10 (oops!)

OpenStudy (anonymous):

57

OpenStudy (phi):

Yes. That is the answer to this long question.

OpenStudy (anonymous):

thats it?

OpenStudy (phi):

Yes. You had to know a lot to solve this. You had to know the sum of the exterior angles add up to 360. You had to add up all the terms and solve for x. You then had to use the x to solve for A.

OpenStudy (anonymous):

Thank you lol (:

OpenStudy (phi):

If you have time, here's a video on solving equations http://www.khanacademy.org/math/algebra/solving-linear-equations/v/solving-equations-2 Khan also has videos on geometry.